A ladder leans against a house with its foot 15 ft from the house. When the foot of the ladder is pulled 9 ft farther away from the house, the top slides down 13 ft. How long is the ladder?
Answer by scott8148(6628)

a^2 + 15^2 = (a - 13)^2 + 24^2

a^2 + 225 = a^2 - 26a + 169 + 576 ___ 26a = 520 ___ a = 20

the two sides of the triangle are 15 and 20
___(looks like a 3-4-5 triangle) so the hypotenuse (ladder) is ?